A steel beam being used in the construction of a skyscraper has a length of 38.000 m when delivered on a cold day at a temperature of 16.000°F. What is the length of the beam when it is being installed later on a warm day when the temperature is 91.000°F? (Give your answer to at least five significant figures.)
Answer #1

First we need to convert Fahrenheit to Celsius

dT = (5/9) x (91 - 16) = 41.67 C

Length, L = Lo (1 + a dT)

L = 38 x (1 + (1.1 x 10^-5) x 41.67) = 38.017 m

Comment in case any doubt please rate my answer....